OpenStudy (anonymous):

Help!!! Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6). Answer choices 7 -14 0 14

OpenStudy (anonymous):

slope = (y2 - y1) / (x2-x1) = (-6 - 8)/(-3+4) = ?

OpenStudy (broskishelleh):

14x + y = -48

OpenStudy (anonymous):

Thanks I have it now :)

OpenStudy (broskishelleh):

So D

OpenStudy (mathstudent55):

@Broskishelleh No, not D.